40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 10 15:41:37 2019, aaron, Update, IOO, WFS measurements Screenshot_from_2019-09-10_18-51-28.pngmctrend_190910.png
    Reply  Wed Sep 11 10:26:56 2019, aaron, Update, IOO, WFS measurements 
       Reply  Wed Sep 11 14:37:43 2019, aaron, Update, IOO, WFS measurements 20190911_WFS.jpg20190911_WFS_2.jpg
          Reply  Thu Sep 12 12:42:31 2019, aaron, Update, IOO, WFS measurements WFS2_TFs.pdfD755499D-9FDF-4E2B-BFC1-016B459DD35D.jpeg
             Reply  Fri Sep 13 10:36:03 2019, aaron, Update, IOO, WFS measurements TF-.pngWFS2_TFs.pdf
                Reply  Mon Sep 16 12:38:59 2019, aaron, Update, IOO, WFS measurements 
                Reply  Thu Oct 3 11:38:35 2019, aaron, Update, IOO, WFS measurements 
          Reply  Mon Sep 30 13:20:55 2019, aaron, Update, IOO, shot noise measurement 
       Reply  Fri Sep 13 10:53:40 2019, aaron, Update, IOO, WFS loop measurements 
          Reply  Mon Sep 16 05:08:04 2019, rana, Update, IOO, WFS loop measurements LSCsensingMatrix.ipynbASCsensingMatrix.ipynb
             Reply  Mon Sep 16 11:55:58 2019, rika, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 09:41:48 2019, gautam, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 10:47:44 2019, rika, Update, IOO, WFS loop measurements MC2.pdf
                   Reply  Wed Sep 18 14:45:52 2019, rika, Update, IOO, WFS loop measurements Screenshot_from_2019-09-18_18-15-34.png
             Reply  Tue Sep 17 10:34:48 2019, aaron, Update, IOO, WFS loop measurements 
                Reply  Tue Oct 8 20:39:42 2019, aaron, Update, IOO, WFS loop measurements 
                   Reply  Wed Oct 9 09:37:28 2019, aaron, Update, IOO, WFS loop measurements 
Message ID: 14875     Entry time: Fri Sep 13 10:36:03 2019     In reply to: 14874     Reply to this: 14882   14929
Author: aaron 
Type: Update 
Category: IOO 
Subject: WFS measurements 

[rika, aaron]

We are at it again. Rika is setting up the TF measurement, I'm looking into scripting the WFS sensing matrix measurement we made earlier in the week so we can return to it next week.

 

Measurement file parameters
WFS2_SEG1 / RFPD

 

 
100 MHz - 500 MHz
WFS2_SEG1 / RFPD   10 MHz - 100 MHz
WFS2_SEG1 / RFPD
 
100 kHz - 10 MHz
WFS2_SEG2 / RFPD TFAG4395A_13-09-2019_181415.txt 100 MHz - 500 MHz
WFS2_SEG2 / RFPD TFAG4395A_13-09-2019_180955.txt 10 MHz - 100 MHz
WFS2_SEG2 / RFPD TFAG4395A_13-09-2019_182918.txt 100 kHz - 10 MHz
WFS2_SEG3 / RFPD TFAG4395A_13-09-2019_121533.txt 100 MHz - 500 MHz
WFS2_SEG3 / RFPD TFAG4395A_13-09-2019_123820.txt 10 MHz - 100 MHz
WFS2_SEG3 / RFPD TFAG4395A_13-09-2019_123243.txt 100 kHz - 10 MHz
WFS2_SEG4 / RFPD
TFAG4395A_13-09-2019_161834.txt
100 MHz - 500 MHz
WFS2_SEG4 / RFPD TFAG4395A_13-09-2019_170007.txt 10 MHz - 100 MHz
WFS2_SEG4 / RFPD TFAG4395A_13-09-2019_172001.txt 100 kHz - 10 MHz

 

When we mesuring TF of SEG4, the beam leaking to SEG1 about 1%.

We finished mesurement SEG2-4 and get the figure by running PDH_calibrate.ipynb .

edit: We observed during segment 2 measurements that blocking the beam reduced the DC level of segment 1 by less than 1%, but still clearly observable. As you can see in the plots, something is suspicious about the normalization of these TFs. We took segment 1 data a few days before the other segments, so perhaps we weren't getting the full beam on the reference PD during the later measurements? When I make this measurement for WFS1, I will try to fix some of these problems by choosing different telescoping optics, and I will consider whether removing the QPD heads from their table will improve the measurement.

Attachment 1: TF-.png  84 kB  Uploaded Fri Sep 13 19:49:27 2019  | Hide | Hide all
TF-.png
Attachment 2: WFS2_TFs.pdf  41 kB  Uploaded Sat Sep 14 11:33:38 2019  | Hide | Hide all
WFS2_TFs.pdf
ELOG V3.1.3-